OpenStudy (anonymous):

Evaluate the integral by using multiple substitutions ∫ [(2r-2)cos√(3(2r-2)^2+9)]/[√(3(2r-2)^2+9)] [dr]

OpenStudy (anonymous):

are you familiar with trigonometric substitution?

OpenStudy (anonymous):

Nope, I'm not good at it. Is this problem similar to the one I asked before?

OpenStudy (anonymous):

i believe so.

OpenStudy (anonymous):

we know that tan(x)^2+1=sec(x)^2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and if \[x=a*tan(\theta) \rightarrow \sqrt{x^2-a^2}=a*sec(\theta)\] this means \[\sqrt{3(2r-2)^2+9}=\sqrt{3}*\sqrt{(2r-2)^2+3}\]

OpenStudy (anonymous):

so we say \[(2r-2)=\sqrt{3}*tan(\theta)\]

OpenStudy (anonymous):

which means we have \[\sqrt{3(2r−2)^2+9}=\sqrt{3}*\sqrt{(2r−2)^2+3}=\sqrt{3}*\sqrt{3}sec(\theta)=3sec(\theta)\]

OpenStudy (anonymous):

does this make sense so far?

OpenStudy (anonymous):

yeah its good. I kinda go lost when you said it was (2r-2)=Sqroot(3)*tan bc i didn't understand where the +9 and +3 went but now i'm good

OpenStudy (anonymous):

*got

OpenStudy (anonymous):

I get it :)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

so now we can substitute our tan and sec functions into the problem.

OpenStudy (anonymous):

\[\frac{(2r-2)cos(\sqrt{(3(2r-2)^2+9)}}{\sqrt{(3(2r-2)^2+9)}}\] becomes \[\frac{\sqrt{3}tan(\theta)cos((3sec(\theta))}{3sec(\theta)}\]

OpenStudy (anonymous):

kk TY I'll post the answer once i'm done with the problem ...^_^

OpenStudy (anonymous):

Ohh XO TY I need to practice these kinds of problems more :/

OpenStudy (anonymous):

we aren't done with this substitution part yet. there is a vital step we still have left.

OpenStudy (anonymous):

you have to solve for dr in terms of d(theta)

OpenStudy (anonymous):

\[(2r−2)=\sqrt{3}tan(θ)\] \[r=\frac{\sqrt{3}tan(θ)+2}{2}\] \[dr=\frac{\sqrt{3}}{2}sec^2(\theta)\]

OpenStudy (anonymous):

whoops i mean \[dr=\frac{\sqrt{3}}{2}sec^2(\theta)d\theta\]

OpenStudy (anonymous):

oh ok lol the theta part went over my head XP

OpenStudy (anonymous):

so \[\int\frac{(2r-2)cos(\sqrt{(3(2r-2)^2+9)}}{\sqrt{(3(2r-2)^2+9}} dr =\int\frac{\sqrt{3}tan(\theta)cos(3sec(\theta))}{3sec(\theta)} *\frac{\sqrt{3}}{2}sec^2(θ)dθ\] \[=\int\frac{sec(\theta)tan(\theta)cos(3sec(\theta))}{2} dθ\]

OpenStudy (anonymous):

well our orinial integral uses dr we need dθ because our funtion is writen in terms of θ

OpenStudy (anonymous):

so if we solve for r in terms of θ and take the derivative we found dr in terms of θ and dθ which is what we needed

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yes I understand. I'll do it over again bc what I had before is all crossed out XP

OpenStudy (anonymous):

Thank you! Didn't think this problem would be so long T.T

OpenStudy (anonymous):

we have a second substitution still...

OpenStudy (anonymous):

we can't integrate cos(3sec(θ))

OpenStudy (anonymous):

we can integrate cos(u) though.

OpenStudy (anonymous):

XO my teacher is nuts. LOL but yes I understand bc it can't stay with theta

OpenStudy (anonymous):

so \[u=3sec(\theta)\] \[du=3sec(\theta)tan(\theta)\]

OpenStudy (anonymous):

i mean \[du=3sec(\theta)tan(\theta)d\theta\]

OpenStudy (anonymous):

so \[\int \frac {sec(θ)tan(θ)cos(3sec(θ))}{2}dθ=\frac{1}{3}\int \frac {cos(3sec(θ))}{2}*3sec(θ)tan(θ)dθ\] \[=\frac{1}{3}\int \frac{cos(u)}{2}du=\frac16sin(u)\]

OpenStudy (anonymous):

\[u=3sec(\theta)\] \[\theta=tan^{-1}(\frac{\sqrt{3}(2r-2)}{3})\]

OpenStudy (anonymous):

\[\frac16sin(u)=\frac16sin(3sec(θ))=\frac16sin(3sec(tan^{-1}(\frac{\sqrt{3}(2r-2)}{3})))\]

OpenStudy (anonymous):

and that is the final answer :) do you get where each part comes from?

OpenStudy (anonymous):

yes I do, but I'm still writing it down and writing side notes to help me know where certain things came from where. Wow its really long thank you so much for your time and patience! I appreciate it =)

OpenStudy (anonymous):

of course :)

OpenStudy (anonymous):

You did an amazing job! You deserve lots of medals and a good rest =)

OpenStudy (anonymous):

thanks :)